OpenStudy (anonymous):

Graph y = ½ x2 and its inverse. List the point(s) of intersection. *Do not use (0, 0)

OpenStudy (anonymous):

x=.5y^2 y^2 = x/(0.5) f^-1(x) = y=sqrt(x/0.5)

OpenStudy (anonymous):

-1, 1) (2, 2) (1.51, 1.51) (-1.25, 1.3)

sam (.sam.):

f(x) = ½ x2 | f^-1 (x)= sqrt(2x) x=-2 y=2 | x=-2 y= No value x=-1 y=1/2 | x=-1 y= No value x=1 y=1/2 | x=1 y= 1.414 x=2 y=2 | x=2 y=2

OpenStudy (anonymous):

o.o

sam (.sam.):

Interception at (2,2) because both have the value

OpenStudy (anonymous):

Thank you Sam! (: